OpenStudy (anonymous):

Help! The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2). Answers: y = 1 over 2x + 3 over 2 y = negative 1 over 2x + 3 over 2 y = 1 over 2x − 3 over 2 y = − 1 over 2x − 3 over 2

OpenStudy (anonymous):

lol guess what step 1 is?

OpenStudy (anonymous):

find the slope?

OpenStudy (anonymous):

right, find the slope of \( y = −2x − 5\) with your eyes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whoops -2

OpenStudy (anonymous):

k good

OpenStudy (anonymous):

step 2 is to know that the perpendicular line will have slope that is the "negative reciprocal" i.e. \(\frac{1}{2}\)

OpenStudy (anonymous):

the slope is going to be negative reciprocal of the given slope which is 1/2

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

lets back up then

OpenStudy (anonymous):

you got the slope of \( y = −2x − 5\) is \(-2\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

questions asks for equation of a line PERPENDICULAR

OpenStudy (anonymous):

that means the slope will be the "negative reciprocal" which is a fancy way of saying "flip it and change the sign"

OpenStudy (anonymous):

i.e for \(-2\) the "negative reciprocal" is \(\frac{1}{2}\)

OpenStudy (anonymous):

so change it to positive

OpenStudy (anonymous):

clear now or no?

OpenStudy (anonymous):

exactly flip it and change the sign

OpenStudy (anonymous):

so 1/2

OpenStudy (anonymous):

if it is negative, make it positive if it is positive, make it negative right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we good now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

substitute the slope 1/2 and the given point (-1,-2) in y-y1=m(x-x1) and solve p.s.m=slope=1/2

OpenStudy (anonymous):

step 3 point is \((−1, −2)\) be careful with the signs point slope formula gives \[y+2=\frac{1}{2}(x+1)\]

OpenStudy (anonymous):

now some fraction arithmetic

OpenStudy (anonymous):

so y=1/2x+2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

pretty close

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

\[y+2=\frac{1}{2}(x+1)\] \[y+2=\frac{1}{2}x+\frac{1}{2}\]

OpenStudy (anonymous):

now subtract \(2\) from both sides

OpenStudy (anonymous):

ok i am waiting is it clear how to remove the parentheses?

OpenStudy (anonymous):

k holding

OpenStudy (anonymous):

y=1/2x - 3/2 is the answer

OpenStudy (anonymous):

y=1/2x-3/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks! can you help me with 2 more?

OpenStudy (anonymous):

you got that negative reciprocal thing now?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yes you change it to positive but hold on let me get the question

OpenStudy (anonymous):

k i'll wait

OpenStudy (anonymous):

Line QR contains (2, 8) and (3, 10) Line ST contains points (0, 6) and (−2, 2). Lines QR and ST are Answers: parallel, because the product of the slopes is −1 perpendicular, because the product of the slopes is −1 parallel, because the slopes are the same perpendicular, because the slopes are the same

OpenStudy (anonymous):

we need to find the slope of both

OpenStudy (anonymous):

find the slopes of both the lines and then compare the formula for finding the slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

slope of \( (2, 8),(3, 10) \) \[m=\frac{10-8}{3-2}=\frac{2}{1}=2\]

OpenStudy (anonymous):

\[(0, 6), (−2, 2)\] let me know what you get

OpenStudy (anonymous):

alright give me a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2

OpenStudy (anonymous):

therefore the answer is C.

OpenStudy (anonymous):

k good

OpenStudy (anonymous):

so the slopes are the same, they are both \(2\)

OpenStudy (anonymous):

so it is c

OpenStudy (anonymous):

that means they are parallel

OpenStudy (anonymous):

yeah, it is always C anyway

OpenStudy (anonymous):

ok last one

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? Answers: 1 over 2 2 negative 1 over 2 −2

OpenStudy (anonymous):

slope is ?

OpenStudy (anonymous):

i mean what is the slope of \( (y−3) = − 2 (x − 4)\) ?

OpenStudy (anonymous):

since it's in point slope dont we need to put it in slope intercept

OpenStudy (anonymous):

oh ya hold on let me solve it

OpenStudy (anonymous):

no you don't need to change the form it is in the point - slope form, all you need is the slope

OpenStudy (anonymous):

slope intercept form

OpenStudy (anonymous):

you see it with your eyes already

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

\[(y−3) = \color{red}{− 2} (x − 4)\]

OpenStudy (anonymous):

its -2 right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now recall that "negative reciprocal" thing

OpenStudy (anonymous):

so postive 2

OpenStudy (anonymous):

flip it

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

"negative reciprocal" has two words, they don't waste words in math yeah \(\frac{1}{2}\)

OpenStudy (anonymous):

so the final answer is 1/2

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

lol 'final answer'

OpenStudy (anonymous):

ok i understand now, thank you!

OpenStudy (anonymous):

who wants to pass a math class?

OpenStudy (anonymous):

yw, hope it is more or less clear